OpenStudy (anonymous):

WILL MEDAL & FAN !!! PLEASE HELP ASAP !! The value of x would be? a. 2 b. 4 c. 8 d. 21

OpenStudy (shinalcantara):

the given equation though

OpenStudy (anonymous):

my bad , this is it .

OpenStudy (gorv):

provide some data ...related to x

OpenStudy (anonymous):

@gorv there was no data , it was just that .

OpenStudy (shinalcantara):

since it is located midway then i presume that the equivalent value would be the average of the two bases.. then we'll have \[4x + 5 = \frac{ 27 + 15 }{ 2 }\] \[4x + 5 = \frac{ 42 }{ 2 }\] \[4x + 5 = 21\] \[4x = 21 - 5\] \[4x = 16\] \[x = 4\]

OpenStudy (anonymous):

Thankyou so much @shinalcantara